OpenStudy (anonymous):

What is the value of the expression? {[24 divided by (6+2)]^4 +1} divided by (2^4 divided by 2) A. 21.5 B. 10.25 C. 5.25 D. 1.25 Please help me someone '.' I wrote the question as best as i could :)

OpenStudy (johnweldon1993):

\[\huge \frac{\frac{24}{(6 + 2)^4} + 1}{\frac{2^4}{2}}\] Like that?

OpenStudy (anonymous):

um, Thats not how it was written out, but it has the right numbers.

OpenStudy (anonymous):

It was just like how I typed it accept when it says 'divided by' those are supposed to be division signs

OpenStudy (johnweldon1993):

\[\huge \frac{(\frac{24}{(6 + 2)})^4 + 1}{\frac{2^4}{2}}\] That? lol I'm trying to figue out how it was written

OpenStudy (johnweldon1993):

Well I'm going to assume that is the correct way since I do receive one of the answer choices available... So we can begin by simplifying the top... what is 24/(6 + 2) .....24 / 8 = ?

OpenStudy (anonymous):

3

OpenStudy (johnweldon1993):

Right...and 3^4 = ? 3 x 3 x 3 x 3 = ?

OpenStudy (anonymous):

81

OpenStudy (johnweldon1993):

And then add the 1....81 + 1 = 82... okay that is the top.. Now the bottom

OpenStudy (johnweldon1993):

What is 2^4? 2 x 2 x 2 x 2?

OpenStudy (anonymous):

16 but then I divided by 2 and got 8 so would i divide 81 and 8?

OpenStudy (johnweldon1993):

82/8 Dont forget that you added the 1 to the 81

OpenStudy (anonymous):

oh yeah I forgot that '.' But I got 10.25 for the final answer, is that right?

OpenStudy (johnweldon1993):

Correct

OpenStudy (anonymous):

THANK YOU! :D

OpenStudy (johnweldon1993):

Anytime :)

OpenStudy (anonymous):

:)